# Long and Short Guidance in Score identity Distillation for One-Step Text-to-Image Generation

### v2

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have shown the capacity to generate photorealistic images consistent with textual descriptions. However, a significant limitation of these models is their slow sample generation, which requires iterative refinement through the same network. In this paper, we enhance Score identity Distillation (SiD) by developing long and short classifier-free guidance (LSG) to efficiently distill pretrained Stable Diffusion models without using real training data. SiD aims to optimize a model-based explicit score matching loss, utilizing a score-identity-based approximation alongside the proposed LSG for practical computation. By training exclusively with fake images synthesized with its one-step generator, SiD equipped with LSG rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Specifically, its data-free distillation of Stable Diffusion 1.5 achieves a record low FID of 8.15 on the COCO-2014 validation set, with a CLIP score of 0.304 at an LSG scale of 1.5, and a FID of 9.56 with a CLIP score of 0.313 at an LSG scale of 2. Our code and distilled one-step text-to-image generators are available at https://github.com/mingyuanzhou/SiD-LSG.
@@ 2.3 Long and Short Guidance
Long strategy: Enhancing CFG of the pretrained score network $f_{\phi}$. Aligning with established practices, the most fundamental approach involves enhancing the CFG of the pretrained score network $f_{\phi}$. An example setting under this strategy in SiD is: $\kappa_{1}=\kappa_{2}=\kappa_{3}=1$ and $\kappa_{4}=3$. The rationale is that by biasing the teacher network $f_{\phi}$ to favor generations more aligned with ${\boldsymbol{c}}$ using a CFG scale of $\kappa_{4}=3$, the student generator is compelled to follow suit. We will present experimental results to demonstrate the effectiveness of this strategy while also discussing its limitations.
@@ 3.1 Long and/or short guidance strategies
Short the CFG of the fake score network. For the short strategy, we set $\kappa_{1}=\kappa_{4}=1$ and explore $\kappa_{2}=\kappa_{3}\in\{0.5,0.125\}$. As illustrated in the right panel of Figure 4, this approach delivers competitive performance, with $\kappa_{2}$ dictating the balance between FID and CLIP scores. However, compared to the long strategy, this configuration generally produces inferior results, as suggested by lower CLIP scores when FIDs are controlled to similar levels. From this figure, it’s also evident that when No CFG is employed, where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1$, the results are underwhelming, underscoring the importance of injecting CFGs into SiD to distill text-to-image diffusion models.
LSG: Long and Short classifier-free Guidance. LSG is an innovative strategy that enhances the performance of SiD in text-guided diffusion distillation. In the simplest LSG configuration, we amplify the CFG during the training of $f_{\psi}$, the fake score network, and reduce it during the training of $G_{\theta}$, the one-step generator. Specifically, to long the CFG of $f_{\psi}$ during its training while shorting it during its evaluation, we set $\kappa_{1}$ as 2 or 3 and set $\kappa_{2}=\kappa_{3}=\kappa_{4}=1$. As depicted in the right panel of Figure 4, setting $\kappa_{1}=3$ (green lines) is as effective as the short strategy with $\kappa_{2}=\kappa_{3}=0.125$ (red lines), and $\kappa_{1}=2$ (black lines) outperforms the short strategy with $\kappa_{2}=\kappa_{3}=0.5$ (blue lines) and is comparable to the long strategy with $\kappa_{4}=3$ (red lines in the left panel). These results validate LSG as an effective guidance strategy to boost SiD’s performance in text-guided diffusion distillation.
Nevertheless, merely matching the best long or short strategy does not suffice to advocate for LSG. However, the performance notably improves if maintaining the CFG scale of the fake score network during the generator’s training after initially increasing it during its own training , effectively balancing FID minimization and CLIP maximization. We explore $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}$ within $\{1.5,2,3,4.5\}$, and present the results in Figure 4 for distilling both SD 1.5 and 2.1-base. Comparisons between SD1.5 outcomes in the left panel of Figure 4 and those in Figure 4 demonstrate superior performance with this LSG setting, indicated by higher CLIP scores at controlled FID levels, and lower FID scores at controlled CLIP levels. Generally, within the range of 1.5 to 4.5, a lower guidance scale correlates with better FID but worse CLIP, and vice versa, as evidenced by the curves for both SD 1.5 and 2.1-base in Figure 4.
@@ 4 Conclusion
In this paper, we introduce a novel method that integrates Classifier-Free Guidance (CFG) with Score identity Distillation (SiD) to efficiently distill Stable Diffusion models into effective one-step generators. Utilizing our innovative long and short CFG strategies (LSG), we have successfully distilled these models using only fake images synthesized by the one-step generator itself. Our approach not only demonstrates the practical potential of SiD, but also sets new benchmarks for one-step diffusion distillation in zero-shot FID scores on the COCO-2014 validation set. The anticipated release of our tools and models is expected to make a significant impact on research and development in generative modeling. This release is designed to enhance efficiency without compromising performance, and it facilitates learning from the teacher model without the need for real images or the addition of extra regression or adversarial losses.

### v3

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have shown the capacity to generate photorealistic images consistent with textual descriptions. However, a significant limitation of these models is their slow sample generation, which requires iterative refinement through the same network. In this paper, we enhance Score identity Distillation (SiD) by developing long and short classifier-free guidance (LSG) to efficiently distill pretrained Stable Diffusion models without using real training data. SiD aims to optimize a model-based explicit score matching loss, utilizing a score-identity-based approximation alongside the proposed LSG for practical computation. By training exclusively with fake images synthesized with its one-step generator, SiD equipped with LSG rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Specifically, its data-free distillation of Stable Diffusion 1.5 achieves a record low FID of 8.15 on the COCO-2014 validation set, with a CLIP score of 0.304 at an LSG scale of 1.5, and an FID of 9.56 with a CLIP score of 0.313 at an LSG scale of 2. Our code and distilled one-step text-to-image generators are available at https://github.com/mingyuanzhou/SiD-LSG.
@@ 2.3 Long and Short Guidance
Long strategy: Enhancing CFG of the pretrained score network $f_{\phi}$. Aligning with established practices, the most common approach involves enhancing the CFG of the pretrained score network $f_{\phi}$. An example setting under this strategy in SiD is: $\kappa_{1}=\kappa_{2}=\kappa_{3}=1$ and $\kappa_{4}=3$. The rationale is that by biasing the teacher $f_{\phi}$ to favor generations more aligned with ${\boldsymbol{c}}$ using a CFG scale greater than 1, such as $\kappa_{4}=3$, the student generator is compelled to follow suit. We will present experimental results to demonstrate the effectiveness of this strategy while also discussing its limitations.
@@ 3.1 Long and/or short guidance strategies
Short the CFG of the fake score network. For the short strategy, we set $\kappa_{1}=\kappa_{4}=1$ and explore $\kappa_{2}=\kappa_{3}\in\{0.5,0.125\}$. As illustrated in the right panel of Figure 4, this approach delivers competitive performance, with $\kappa_{2}$ dictating the balance between FID and CLIP scores. However, compared to the long strategy, this configuration generally produces inferior results, as suggested by lower CLIP scores when FIDs are controlled to similar levels. From Figure 4, it’s also evident that when No CFG is employed, where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1$, the results are underwhelming, underscoring the importance of injecting CFGs into SiD to distill text-to-image diffusion models.
LSG: Long and Short classifier-free Guidance. Below, we explore how to effectively integrate the long and short strategies to enhance text guidance in diffusion distillation. Initially, we discovered an alternative approach to shorting the CFG of the fake score network $f_{\psi}$: by amplifying the CFG during the training of $f_{\psi}$. This approach represents the “simplest” form of LSG strategy. Specifically, we set $\kappa_{1}$ to 2 or 3, while keeping $\kappa_{2}=\kappa_{3}=\kappa_{4}$ at 1. As illustrated in the right panel of Figure 4, setting $\kappa_{1}=3$ (green lines) proves as effective as the short strategy with $\kappa_{2}=\kappa_{3}=0.125$ (red lines), and $\kappa_{1}=2$ (black lines) outperforms the short strategy with $\kappa_{2}=\kappa_{3}=0.5$ (blue lines) and is comparable to the long strategy with $\kappa_{4}=3$ (red lines in the left panel). These findings validate this “simplest” LSG as an effective guidance strategy in text-guided diffusion distillation.
Nevertheless, merely matching the performance of the best long or short strategy is insufficient to justify adopting this “simplest” LSG. The efficacy of LSG is notably enhanced when the CFG scale applied to the fake score network during its training exceeds 1 and is maintained throughout the generator’s training. This strategy strikes an effective balance between minimizing FID and maximizing CLIP. Specifically, for this recommended LSG configuration, we evaluated $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}$ values within $\{1.5,2,3,4.5\}$, presenting the results in Figure 4 for distilling both SD 1.5 and 2.1-base models. Comparisons between SD1.5 outcomes in the left panel of Figure 4 and those in Figure 4 demonstrate superior performance with this LSG setting, indicated by higher CLIP scores at controlled FID levels, and lower FID scores at controlled CLIP levels. Generally, within the range of 1.5 to 4.5, a lower guidance scale correlates with better FID but worse CLIP, and vice versa, as evidenced by the curves for both SD 1.5 and 2.1-base in Figure 4.
@@ 3.3 Quantitative and qualitative evaluations
SD1.5 distilled with SiD-LSG ($\kappa=1.5$): FID $=$ 8.15, CLIP $=$ 0.304 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
SD2.1-base distilled with SiD-LSG ($\kappa=4.5$): FID = 16.54, CLIP = 0.322 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Figure 7: Visual comparison of two SiD-LSG models: one preferred for FID and the other for CLIP. All images are generated from the same text prompt: “A distinguished older gentleman in a vintage study, surrounded by books and dim lighting, his face marked by wisdom and time. 8K, hyper-realistic, cinematic, post-production." The model with a lower guidance scale of $\kappa=1.5$, which achieves a record-low one-step-generation FID of 8.15 and a competitive CLIP score of 0.304, produces images that are more diverse but align less closely with specific text details, such as “dim lighting.” Conversely, the model with a higher guidance scale of $\kappa=4.5$, achieving a high CLIP score of 0.322 and noted for state-of-the-art human preference scores (HPSv2) as shown in Table 2, presents a relatively high FID of 16.54, indicating less diversity but superior text alignment and visual quality.
SD1.5 distilled with SiD-LSG ($\kappa=1.5$): FID $=$ 8.15, CLIP $=$ 0.304 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
SD2.1-base distilled with SiD-LSG ($\kappa=4.5$): FID = 16.54, CLIP = 0.322 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Figure 8: Analogous to Figure 8, this plot compares two rows of images generated using two distinct guidance scales, both conditioning on the same text prompt: “poster art for the collection of the asian woman, in the style of gloomy, dark orange and white, dynamic anime, realistic watercolors, nintencore, weathercore, mysterious realism –ar 69:128 –s 750 –niji 5".
@@
To contextualize the numerical differences in metrics such as FID, CLIP, and HPSv2, we present Figures 8 and 8 to illustrate what these differences mean for visual perception: the model with the best FID score excels in diversity, while the model with the best CLIP score stands out in text alignment and aesthetic quality.
@@ 4 Conclusion
In this paper, we introduce a novel approach that combines Classifier-Free Guidance (CFG) with Score identity Distillation (SiD) to efficiently distill Stable Diffusion models into effective one-step generators. Utilizing our innovative Long and Short CFG strategies (LSG), we have distilled these models using only synthetic images generated by the one-step generator. This method not only validates the practical potential of SiD but also establishes new benchmarks for one-step diffusion distillation, achieving remarkable zero-shot FID scores on the COCO-2014 validation set. Designed for enhanced efficiency without sacrificing performance, our approach allows for learning from the teacher model without requiring real images or incorporating additional regression or adversarial losses. We have made our code and distilled models publicly available to facilitate further research.
@@ Appendix B Related work
How does SiD-LSG differ from previous methods? SiD-LSG introduces several unique features that distinguish it from previous text-to-image diffusion distillation methods. Firstly, similar to SwiftBrush [Nguyen and Tran, 2024], SiD-LSG is a data-free distillation method, which eliminates the need for original training datasets or synthetic data generated with SDE/ODE solvers during the distillation process. However, unlike SwiftBrush, SiD-LSG necessitates gradient backpropagation through the score networks, a critical step akin to gradient backpropagation through the discriminator in methods employing adversarial losses. Additionally, SiD-LSG applies CFG to both the training and evaluation of the fake score network. This is a departure from previous methods, which typically apply CFG only during the evaluation of the pretrained score network. Finally, SiD-LSG aims to minimize a model-based explicit score-matching loss, a type of Fisher divergence. In contrast, previous methods often focus on minimizing losses based on KL divergence, consistency, regression, GAN-based adversarial tactics, or a combination thereof.